

********************************************************************
UARS WEEKLY STATUS REPORT - 6 SEP 2002
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
Through 9/2/02 when it was turned off for the yaw-around, SUSIM
successfully performed its daily mid-resolution and MgII
solar scans during periods of predicted UARS telemetry.
Reprocessing of SUSIM data for the period 10/12/91 through 4/27/02
to version V21 is complete. Irradiance data are available from the
SUSIM ftp site. A new version (V21r2) of the MgII core-to-wing
ratio also may be found there.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
PEM continues to work with the GUVI team on the diagnosis of data from
the first conjunction period. Analysis of the two conjunction passes showed
that the aurora as detected by GUVI was at higher latitude than UARS.
GUVI images from other times from this conjunction period show spatial
regions that both GUVI and PEM sampled when aurora was present. Separation
is several minutes to 10's of minutes.
PEM data from the second GUVI-PEM conjunction has been processed and is
now being analyzed.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in the UARS Monthly Status Review
Meeting held on 5 September 2002.
The MPG continues to serve as the point-of-contact for UARS
instrument operations to support re-engineering of the UARS
Mission Operations Concept to develop capabilities for the
Science Traceability Mission.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 24 August 2002 (Orbit 59888, GMT Day 236) through
30 August 2002 (Orbit 59992, GMT Day 242).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HRDI 238/0058 59919 DAY ONLY MODE
HALOE 242/0156 59979 POWER OFF
The beta angle ranged between the angles of 38.1 degrees to 20.5
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 04 September 2002 (DOY 247).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.92 to 1.91 degrees C. Additional
battery performance data for each battery on 23 August and 30 August
2002 are provided in Appendix D.
The clock error ranged from -7.5 to 1.3 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
HRDI = 01
HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 01
Four (4) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The Raytheon development team has delivered a new software delivery,
which takes care of several minor problems that exist on the current
Eclipse system. This patch has yet to be tested on the prototype
system due to a hard disk failure, but should be tested within the
next couple weeks.
MEETINGS
- The FOT attended an ERBS transition meeting on 27 August 2002.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 04 September 2002 (DOY 247).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 30 August 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
59919 238/0058 02-304 HRDI DAY ONLY MODE
59979 242/0156 02-305 HALOE POWER OFF
_____________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 30 August 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
______________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 30 August 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
-None this report period.
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 30 August 2002
23 August 2002 - GMT Day 235 Beta = 40.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -134.4/-156.8 -2.73/-3.68 +00.0/+00.0 08.5 0.000 0.0
2 +11.2/-67.2 +6.11/+4.51 +27.2/-16.0 23.7 1.030 21.1
3 +33.6/-33.6 +5.15/+3.87 +26.8/-16.0 23.7 1.030 20.7
30 August 2002 - GMT Day 242 Beta = 20.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -134.4/-156.8 -3.05/-4.62 +00.0/+00.0 08.5 0.000 0.0
2 +11.2/-44.8 +5.47/+3.55 +28.0/-13.6 24.2 1.040 19.1
3 +22.4/-39.2 +4.51/+2.60 +26.8/-13.2 24.2 1.030 18.4
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 30 August 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-048 FOT N/A 242/1300 N/A Loss of UPS-CMS connectivity
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-045 FOT 59717 224/1458 14170 Generic Late Acquisition
02-046 FOT 59831 232/0534 30824 Data dropout due to an
equipment fail-over at WSGT.
02-047 SSPP 59844 233/0213 N/A SSPP in wrong mode
02-048 FOT N/A 242/1300 N/A Loss of UPS-CMS connectivity